# Arithmetic on Continued Fractions

### v1

@@
Abstract.
Gosper developed algorithms for adding, subtracting, multiplying, or dividing two CFs, and for solving quadratics with CF coefficients, getting a CF as the result. Here we present modified versions of those algorithms which avoid all difficulties with infinite loops. We have implemented these algorithms in Haskell.
1. Introduction
A continued fraction is a (possibly infinite) expression of the form
@@ 1. Introduction
where the terms $a_{i}$ are integers, with $a_{i}$ positive when $i>0$. It is written more compactly as $[a_{0},a_{1},\cdots]$. The properties of continued fractions are very well-known [^1] https://en.wikipedia.org/wiki/Simple_continued_fraction, so here we only remind the reader of some notation we will use:
  • A quadratic irrational has a periodic CF expansion. We denote the periodic part by putting it in parentheses, i.e $[1,(2,3)]=[1,2,3,2,3,2,3,\cdots]$.
  • A rational number has a finite CF expansion $[a_{0},a_{1},\cdots a_{k}]$. It will be convenient to consider a finite CF as ending with an infinite term $a_{k+1}=\infty$, treating $\frac{1}{\infty}$ as equal to zero.
Gosper (in an appendix to the famous HAKMEM report [1]) developed algorithms for adding, subtracting, multiplying, or dividing two CFs, and for solving quadratics with CF coefficients, getting a CF as the result; the point of course is that we can do this entirely within the CF representation, making no use of floating-point arithmetic. Here we present modified versions of those algorithms which avoid all difficulties with infinite loops; any combination of arithmetic computations can be carried out to any required degree of accuracy in a finite number of steps. We have implemented these algorithms in Haskell [^2] https://github.com/mjcollins10/ContinuedFractions.
2. Arithmetic on One CF
Before we describe how to add or multiply two CFs, we consider the simpler problem of operations combining a single CF with a rational number $p/q$. A few examples reveal there is no evident general pattern for transforming the CF terms of irrational $x$ into the terms of $px/q$ or $x+p/q$, even in the apparently simplest cases:
@@ 2. Arithmetic on One CF
[…]
It will be convenient to have notation for the “tail” of a continued fraction, so let
@@
$[y_{1},y_{2},\cdots]=1+\frac{r_{0}-1}{r_{0}+1}=y_{1}+\frac{1}{[y_{2},y_{3}\cdots]}$
$[y_{2},\cdots]=\frac{r_{0}+1}{r_{0}-1}\ .$
Now the floor of $\frac{r_{0}+1}{r_{0}-1}$ ranges from 1 to infinity as $r_{0}$ ranges from one to infinity, so we need to make use of $\pi_{1}=7$; we substitute $r_{0}=7+1/r_{1}$ to get
$[y_{2},\cdots]=\frac{8+1/r_{1}}{6+1/r_{1}}=\frac{8r_{1}+1}{6r_{1}+1}=1+\frac{2r_{1}}{6r_{1}+1}=1+\frac{1}{[y_{3},\cdots]}\ .$
@@
$[y_{3},\cdots]=\frac{6r_{1}+1}{2r_{1}}=3+\frac{1}{2r_{1}}=3+\frac{1}{[y_{4},\cdots]}\ .$
Going further will require substituting $r_{1}=15+1/r_{2}$ into $2r_{1}=[y_{4},\cdots]$. It is now clear that we will generate terms of $y$ by repeatedly determining the integer part (i.e. floor) of expressions of the form
(1) $\frac{px+q}{rx+s}$
where $x>1$ and the continued fraction expansion of $x$ is known. Functions of this form are called homographic. We will identify a homographic function with the matrix $\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$.
To formalize our observations, we first define the range of the function given by (1) as the set of possible floors, i.e.
[…]
We will be able to produce the next term of output when this set is a singleton. If $-s/r\leq 1$ then the denominator cannot be zero, and the upper and lower limits are the min and max of $\{p/r,(p+q)/(r+s)\}$: if $p/r$ is the max, and is also an integer, then the integer part cannot be larger than $p/r-1$, since we are approaching $p/r$ from below. If $-s/r>1$ the range will be infinite, so there is no need to compute it explicitly; we need to transform the expression by incorporating more information about $x$.
We define two transformations on homographics. When the range of the homographic expression for $[y_{j},y_{j+1},\cdots]$ is not a singleton set, we do not yet have enough information to determine $y_{j}$, and must ingest the next term of $x$. If this term is $k$ we make the substitution $x\leftarrow k+1/x$, leading to the definition
(3) $\mbox{ingest}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))=\begin{pmatrix}kp+q&p\\ kr+s&r\end{pmatrix}$
When the range of the homographic expression for $[y_{j},y_{j+1},\cdots]$ consists of a single integer $k$, we can produce the next output term $y_{j}=k$; the rest of the output, $[y_{j+1},y_{j+2},\cdots]$, is the continued fraction expansion of
@@
(4) $\mbox{produce}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))=\begin{pmatrix}r&s\\ p-kr&q-ks\end{pmatrix}\ .$
Note that if we started with rational $x$, we will eventually reach $x_{i}=\infty$; ingesting infinity returns the limit ${\left(\begin{smallmatrix}0&p\\ 0&r\end{smallmatrix}\right)}$. This is the constant function at the rational number $p/q$, so there will never be any need to ingest the (nonexistent) subsequent terms of $x$. If we proceed as in the examples above, we will compute the CF representation of $p/q$ as the last part of the finite expansion of $y$, ending with a production that subtracts the final term $y_{k}$ from the expression $\frac{y_{k}}{1}$; this leads to $M=\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)$, which we may read as infinity.
@@
  $M\leftarrow\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$the initial matrix, i.e. the expression whose CF we will compute}
  $M\leftarrow\mbox{ingest}(x_{0},M)$
@@
  while $M\neq\infty$ do
     while $\mathcal{R}(M)$ is not a singleton do
@@
     $y_{j}\leftarrow\mathcal{R}(M)$
@@
Figure 1. Algorithm for arithmetic on a CF and a rational number
@@
So now we have an algorithm for producing the CF expansion $[y_{0},y_{1},\cdots]$ of $\frac{px+q}{rx+s}$, assuming the expansion of $x$ is known. To compute $x/k$ we would start with $M=\left(\begin{smallmatrix}1&0\\ 0&k\end{smallmatrix}\right)$; to get $kx$ we would start with $M=\left(\begin{smallmatrix}k&0\\ 0&1\end{smallmatrix}\right)$; while $x+\frac{j}{k}$ would be $M=\left(\begin{smallmatrix}k&j\\ 0&k\end{smallmatrix}\right)$.
2.1. Properties of the Algorithm
@@
The outer loop “while $M\neq\infty$” is an infinite loop if $\mathbf{x}=[\mathbf{x}_{0},\mathbf{x}_{1},\cdots]$ is irrational (here we use boldface $\mathbf{x}$ for the real-valued input, as distinct from the symbolic variable $x$ in the homomorphic expression). This can be implemented quite directly in a language like Haskell333haskell.org with lazy evaluation; such languages support conceptually infinite data structures, which are a perfect fit for working with continued fractions. Under lazy evaluation, defining a variable $y$ to be the result of an algorithm just associates $y$ with the finite recursive expression that defines the algorithm; the number of $y$ terms actually generated will be only what is needed by subsequent computations.
@@
The inner loop “while $\mathcal{R}(M)$ is not a singleton” is guaranteed to terminate; we will never be in an infinite loop waiting for the next term. We have already observed that this loop terminates if $\mathbf{x}$ is rational, so assume it is irrational. As we iterate through this loop, we are repeatedly rewriting the initial $M=\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$ with $[\mathbf{x}_{0},\cdots\mathbf{x}_{i},x]$ in place of $x$. As $i$ goes to infinity, the upper and lower bounds on $[\mathbf{x}_{0},\cdots\mathbf{x}_{i},x]$ approach $\mathbf{x}$, so the upper and lower bounds of the matrix approach $\frac{p\mathbf{x}+q}{r\mathbf{x}+s}$. This limit is irrational, so eventually the bounding interval contains no integers; at this point $\mathcal{R}(M)$ will be a singleton (namely the common floor of all numbers in the interval), and we will produce a term.
2.2. Visualizing the Algorithm
@@
Since $\mbox{ingest}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))$ returns a matrix whose right column is $\left(\begin{smallmatrix}p\\ r\end{smallmatrix}\right)$, and $\mbox{produce}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))$ returns a matrix whose top row is $(r\ s)$, we can visualize the progress of the algorithm as moving a $2\times 2$ window through a two-dimensional grid of integers. We start with $M$ in the upper right corner; ingest moves left and produce moves down. We put the $x_{i}$ on the top row, at the positions where we ingest, and $y_{j}$ in the right column, at the positions where we produce. Here is such a representation of our earlier example of $\pi/2$, taken further to use $\pi=[3,7,15,1\cdots]$ to obtain $\pi/2=[1,1,1,3,31,\cdots]$.
$\begin{matrix}&&&1&15&7&3&&\\ &&&&&3&1&0&\\ &&&&&2&0&2&1\\ &&&&8&1&1&&1\\ &&&&6&1&-1&&1\\ &&32&30&2&0&&&3\\ &&1&1&0&1&&&31\\ &&1&-1&&&&&\end{matrix}$
3. Arithmetic on Two Continued Fractions
@@
[…]
with $[y_{0},y_{1},\cdots]=\sqrt{2}=[1,(2)]$. We can write this sum as
@@
Substituting $[\pi_{1},\cdots]=7+1/[\pi_{2},\cdots]$ and $[y_{1},\cdots]=2+1/[y_{2},\cdots]$ leads (after a lot of high-school algebra) to
@@
The integer parts of $\frac{65}{14},\frac{29}{7},\frac{9}{2}$ and $\frac{4}{1}$ are all 4; therefore $\lfloor z\rfloor=4$. The next term will be the floor of
@@
into which we can substitute $[\pi_{2},\cdots]=15+\frac{1}{[\pi_{3},\cdots]},[y_{2},\cdots]=2+\frac{1}{[y_{3},\cdots]}$ and so on.
[…]
$(3+\frac{1}{[\pi_{1},\cdots]})(1+\frac{1}{[y_{1}\cdots]})=\frac{3[\pi_{1},\cdots][y_{1}\cdots]+3[\pi_{1},\cdots]+[y_{1}\cdots]+1}{[\pi_{1},\cdots][y_{1}\cdots]}\ .$
@@
$\left\{\frac{a}{e},\frac{a+b}{e+f},\frac{a+c}{e+g},\frac{a+b+c+d}{e+f+g+h}\right\}\ .$
@@
(6) $\mbox{produce}(k,\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right))=\left(\begin{smallmatrix}e&f&g&h\\ (a-ke)&(b-kf)&(c-kg)&(d-kh)\end{smallmatrix}\right)$
If the floor is not determined, we must narrow its range by ingesting the next term of either $x$ or $y$. If we use $y=[s,y_{k+1},y_{k+2}\cdots]$, we make the substitution $y\leftarrow s+1/y$ to get
$\frac{(sa+b)xy+ax+(sc+d)y+c}{(se+f)xy+ex+(sg+h)y+g}$
@@
[…]
@@
[…]
for the result of substituting $x\leftarrow s+1/x$.
@@
  $M\leftarrow\mbox{ingest\_y}(y_{0},\mbox{ingest\_x}(x_{0},M))$
@@
Figure 2. Preliminary algorithm for arithmetic on two CFs
We now have a preliminary algorithm for arithmetic; this will in fact require substantial modification, but we include it as a summary of the key ideas. It is not much different from the single-CF case. Computation of $x+y,x-y,xy$, and $x/y$ begin with matrices $\left(\begin{smallmatrix}0&1&1&0\\ 0&0&0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}0&1&-1&0\\ 0&0&0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}1&0&0&0\\ 0&0&0&1\end{smallmatrix}\right)$ and $\left(\begin{smallmatrix}0&1&0&0\\ 0&0&1&0\end{smallmatrix}\right)$.
@@
3.1. Failure to Converge
Algorithm (2) demonstrates the basic idea of CF arithmetic, but is insufficient because it can fail to converge; we might endlessly ingest terms of both $x$ and $y$ without ever obtaining a bihomographic expression whose floor is known. The simplest case is multiplying $\sqrt{2}$ by itself. The first iteration, ingesting 1 from $x$ and from $y$, gives
$z=\sqrt{2}*\sqrt{2}=[z_{0},z_{1},\cdots]={\frac{1xy+1x+1y+1}{1xy+0x+0y+0}}\ .$

### v2

# Transcendental Functions on Continued Fractions

@@
Abstract
Gosper developed an algorithm for performing arithmetic operations on continued fractions (CFs), getting a CF as the result. Straightforward implementation of the algorithm leads to infinite loops on some inputs. Here we present a modified version of the arithmetic algorithm and prove that it avoids all difficulties with infinite loops. We then combine CF arithmetic with the “spigot” algorithm to compute the CF expansions of exponential, logarithmic, and trigonometric functions of CFs. We have implemented these algorithms in Haskell.
1 Introduction
A simple continued fraction [^1] https://en.wikipedia.org/wiki/Simple_continued_fraction is a (possibly infinite) expression of the form
@@ 1 Introduction
where the terms $a_{i}$ are integers. It can be written more compactly as $a_{0},a_{1},\cdots]$ [^2] Many authors write this as $a_{0};a_{1},\cdots]$. Here we consider only regular simple continued fractions, which means $a_{i}\geq 1$ when $i>0$. The properties of continued fractions, and the motivations for studying them, are very well-known [12], so here we only remind the reader of some notation we will use:
  • A quadratic irrational has a periodic CF expansion. We denote the periodic part by putting it in parentheses, i.e $1,(2,3)]=[1,2,3,2,3,2,3,\cdots]$.
  • A rational number has a finite CF expansion $a_{0},a_{1},\cdots a_{k}]$. It will be convenient to consider a finite CF as ending with an infinite term $a_{k+1}=\infty$, equating $\frac{1}{\infty}$ with zero.
Gosper (in an appendix to the famous HAKMEM report [1]) developed an algorithm for performing arithmetic on two CFs, getting a CF as the result; the point of course is that we can do this entirely within the CF representation, making no use of floating-point operations. Straightforward implementation of the algorithm leads to infinite loops on some inputs [^3] https://srossd.com/blog/2020/gosper-1/ (section 3.1). Here we present a modified version of the arithmetic algorithm, and prove that these modifications avoid infinite loops in the following sense: given any $\varepsilon>0$, if the correct result is $z=[z_{0},z_{1},\cdots]$, our algorithm will eventually produce a finite CF $z_{0},\cdots z_{k-1},z^{\prime}_{k}]$ with proof that
$|z-[z_{0},\cdots z_{k-1},z^{\prime}_{k}]|<\varepsilon\ .$
We then extend the arithmetic algorithm with the techniques of the spigot algorithm for the digits of $\pi$ [3, 11], to compute CF representations of $e^{x},\log(x)$, $\cos(x)$, and $\arcsin(x)$, where $x$ is given as a CF. Of course, since we can do arithmetic, we could compute any $n^{th}$ degree Taylor approximation of such functions to any degree of accuracy, but we go further; we develop an iterative algorithm that continually generates terms of the CF, using only as much of the series (and as many terms of $x$) as are required to determine the next output term.
We have implemented these algorithms in Haskell [^4] https://github.com/mjcollins10/ContinuedFractions.
2 Arithmetic on One CF
Before we describe how to add or multiply two CFs, we consider the simpler problem of operations combining a single CF with a rational number $p/q$. It is not in fact necessary to handle this as a special case, but considering this case separately allows us to present the core idea of the algorithm in a straightforward way.
A few examples reveal there is no evident general pattern for transforming the CF terms of irrational $x$ into the terms of $px/q$ or $x+p/q$, even in the apparently simplest cases:
@@ 2 Arithmetic on One CF
$y_{0},y_{1},y_{2}\cdots]=\frac{\pi}{2}=1.5707963267948966\cdots\ .$
Let $\pi=[\pi_{0},\pi_{1},\cdots]$ (in general we use $x_{i}$ for the $i^{\mbox{th}}$ CF term of a real number $x$). It will be convenient to have notation for the “tail” of a continued fraction, so let
@@
$y_{1},y_{2},\cdots]=1+\frac{r_{0}-1}{r_{0}+1}=y_{1}+\frac{1}{[y_{2},y_{3}\cdots]}$
$y_{2},\cdots]=\frac{r_{0}+1}{r_{0}-1}\ .$
Now the floor of $\frac{r_{0}+1}{r_{0}-1}$ ranges from one to infinity as $r_{0}$ ranges from one to infinity, so we need to make use of $\pi_{1}=7$; we substitute $r_{0}=7+1/r_{1}$ to get
$y_{2},\cdots]=\frac{8+1/r_{1}}{6+1/r_{1}}=\frac{8r_{1}+1}{6r_{1}+1}=1+\frac{2r_{1}}{6r_{1}+1}=1+\frac{1}{[y_{3},\cdots]}\ .$
@@
$y_{3},\cdots]=\frac{6r_{1}+1}{2r_{1}}=3+\frac{1}{2r_{1}}=3+\frac{1}{[y_{4},\cdots]}\ .$
Going further will require substituting $r_{1}=15+1/r_{2}$ into $y_{4},\cdots]=2r_{1}$. It is now clear that we will generate terms of $y$ by repeatedly determining the integer part (i.e. floor) of expressions of the form
$\frac{px+q}{rx+s}$ (1)
where $x>1$, the coefficients are integral, and the continued fraction expansion of $x$ is known. Functions of this form are called homographic. We will identify a homographic function with the matrix $M=\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$ and may write (1) as $M(x)$. We can assume without loss of generality that the determinant adbcad-bc is nonzero, and $\gcd(a,b,c,d)=1$.
If $s/r\leq 1$ then the denominator cannot be zero, and $M(x)$ is between the min and max of $\{p/r,(p+q)/(r+s)\}$; if
$\left\lfloor\frac{p}{r}\right\rfloor=\left\lfloor\frac{p+q}{r+s}\right\rfloor$
then this common value is $\lfloor M(x)\rfloor$; we say that $M$ is unambiguous.
We always have
$M([x_{i},x_{i+1},\cdots])=[y_{j},y_{j+1},\cdots]$ (2)
where $x_{i},x_{i+1},\cdots]$ is the unread part of the input and $y_{0},\cdots y_{j-1}]$ is the output prefix already computed. We define two transformations on $M$ which preserve (2). When $M$ is ambiguous, we do not yet have enough information to determine $y_{j}$, and must ingest the next term of $x$. If this term is $x_{i}=k$ we make the substitution $x\leftarrow k+1/x$, leading to the definition
$\mbox{ingest}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))=\begin{pmatrix}q+kp&p\\ s+kr&r\end{pmatrix}\$.
When $M$ is unambiguous with $k=\left\lfloor\frac{p}{r}\right\rfloor$, we can produce the next output term $y_{j}=k$; the rest of the output, $y_{j+1},y_{j+2},\cdots]$, is the continued fraction expansion of
@@
$\mbox{produce}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))=\begin{pmatrix}r&s\\ p-kr&q-ks\end{pmatrix}\$.
Note that these two transformations do not change the absolute value of the determinant of $M$.
If we started with rational $x$, we will eventually reach $x_{i}=\infty$; ingesting infinity returns the limit $\left(\begin{smallmatrix}0&p\\ 0&r\end{smallmatrix}\right)$}. This is the constant function at the rational number $p/r$, so there will never be any need to ingest the (nonexistent) subsequent terms of $x$. If we proceed as in the examples above, we will compute the CF representation of $p/r$ as the last part of the finite expansion of $y$, ending with a production that subtracts the final term $y_{k}$ from the expression $\frac{y_{k}}{1}$; this leads to $M=\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)$, which we may read as infinity.
 {input $x=[x_{0},x_{1},\cdots]$}
 {input $M=\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$}
 {output $y$: CF expansion of M(x)M(x)}
@@
$M\leftarrow\mbox{ingest}(x_{0},M)$ {updating $M$ updates variables p,q,r,sp,q,r,s}
@@
while $M\neq\left(\begin{smallmatrix}0&1\\ 0&0\end{smallmatrix}\right)$ do
  while $M$ is ambiguous do
@@
  $y_{j}\leftarrow\left\lfloor\frac{p}{r}\right\rfloor$
@@
Figure 1: Algorithm for arithmetic on a CF and a rational number
@@
So now we have an algorithm for producing the CF expansion $y_{0},y_{1},\cdots]$ of $\frac{px+q}{rx+s}$, assuming the expansion of $x$ is known. To compute $x/k$ we would start with $M=\left(\begin{smallmatrix}1&0\\ 0&k\end{smallmatrix}\right)$; to get $kx$ we would start with $M=\left(\begin{smallmatrix}k&0\\ 0&1\end{smallmatrix}\right)$; while $x+\frac{j}{k}$ would be $M=\left(\begin{smallmatrix}k&j\\ 0&k\end{smallmatrix}\right)$. Note that computing any homomorphic expression $\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$ is no more difficult than a single arithmetic operation.
2.1 Proof of Termination
@@
The outer loop “while $M\neq\infty$” is an infinite loop if $\mathbf{x}=[\mathbf{x}_{0},\mathbf{x}_{1},\cdots]$ is irrational (here we use boldface $\mathbf{x}$ for the real-valued input, as distinct from the symbolic variable $x$ in the homographic expression). This loop can be implemented quite directly in a language like Haskell555haskell.org with lazy evaluation[5]; such languages support conceptually infinite data structures, which are a perfect fit for working with continued fractions. Under lazy evaluation, defining a list-type variable $y$ to be the result of an algorithm just associates $y$ with the finite recursive expression that defines the algorithm; the number of elements of $y$ actually generated will be only what is needed by subsequent computations.
@@
The inner loop “while $M$ is ambiguous” is guaranteed to terminate even on irrational $\mathbf{x}$. As we iterate through the inner loop, we are repeatedly rewriting the initial $M=\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right)$ with $\mathbf{x}_{0},\cdots\mathbf{x}_{i},x]$ in place of $x$. The upper and lower bounds on $\mathbf{x}_{0},\cdots\mathbf{x}_{i},x]$ approach $\mathbf{x}$, so the upper and lower bounds of the matrix approach $\frac{p\mathbf{x}+q}{r\mathbf{x}+s}$. Since this limit is irrational666because we have ruled out the degenerate case of zero determinant, eventually the bounding interval contains no integers; at this point $M$ will be unambiguous and we will produce a term.
Liardet and Stambul [7] go beyond this basic observation to prove a bound (dependent on the determinant of $M$) on the maximum number of iterations needed to get the next term. Miska, Murru, and Romeo [9] give performance results suggesting that, almost always, four input terms are sufficient to get the next output.
2.2 Visualizing the Algorithm
@@
Since $\mbox{ingest}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))$ returns a matrix whose right column is $\left(\begin{smallmatrix}p\\ r\end{smallmatrix}\right)$, and $\mbox{produce}(k,\left(\begin{smallmatrix}p&q\\ r&s\end{smallmatrix}\right))$ returns a matrix whose top row is (rs)(r\ s), we can visualize the progress of the algorithm as moving a $2\times 2$ window through a two-dimensional grid of integers. We start with $M$ in the upper right corner; ingest moves left and produce moves down. We put the $x_{i}$ on the top row, at the positions where we ingest, and $y_{j}$ in the right column, at the positions where we produce. Here is such a representation of our earlier example of $\pi/2$, taken further to use $\pi=[3,7,15,1\cdots]$ to obtain $\pi/2=[1,1,1,3,31,\cdots]$.
[…]
3 Arithmetic on Two Continued Fractions
@@
$z=\pi+\sqrt{2}=4.555806215962888\cdots$
with $y_{0},y_{1},\cdots]=\sqrt{2}=[1,(2)]$. We can write this sum as
@@
Substituting $\pi_{1},\cdots]=7+1/[\pi_{2},\cdots]$ and $y_{1},\cdots]=2+1/[y_{2},\cdots]$ leads (after a lot of high-school algebra) to
@@
The integer parts of $\frac{65}{14},\frac{29}{7},\frac{9}{2}$ and $\frac{4}{1}$ are all 4; therefore $z_{0}=\lfloor z\rfloor=4$. This conclusion follows easily from the mediant inequality: if $\frac{a}{c}<\frac{b}{d}$ with all terms positive, then
$\frac{a}{c}<\frac{a+b}{c+d}<\frac{b}{d}$
The next term $z_{1}$ will be the floor of
@@
into which we can substitute $\pi_{2},\cdots]=15+\frac{1}{[\pi_{3},\cdots]},[y_{2},\cdots]=2+\frac{1}{[y_{3},\cdots]}$ and so on.
Similarly, the product $\pi\sqrt{2}=4.442882938158366\cdots$ is
$(3+\frac{1}{[\pi_{1},\cdots]})(1+\frac{1}{[y_{1}\cdots]})=\frac{3[\pi_{1},\cdots][y_{1}\cdots]+3[\pi_{1},\cdots]+[y_{1}\cdots]+1}{[\pi_{1},\cdots][y_{1}\cdots]}$
@@
$\left\{\frac{a}{e},\frac{a+b}{e+f},\frac{a+c}{e+g},\frac{a+b+c+d}{e+f+g+h}\right\}$
where we may ignore fractions with numerator and denominator both zero.
@@
$\mbox{produce}(k,\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right))=\left(\begin{smallmatrix}e&f&g&h\\ (a-ke)&(b-kf)&(c-kg)&(d-kh)\end{smallmatrix}\right)$
If the floor is not determined (in particular when the denominator might be zero), we must narrow the range by ingesting the next term of either $x$ or $y$. If we use $x=[s,x_{k+1},x_{k+2}\cdots]$, we make the substitution $x\leftarrow s+1/x$ to get
$\frac{(sa+c)xy+(sb+d)x+ay+b}{(ae+g)xy+(af+h)x+ey+f}$
@@
$\mbox{ingest\_x}(s,\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right))=\left(\begin{smallmatrix}(sa+c)&(sb+d)&a&b\\ (ae+g)&(af+h)&e&f\end{smallmatrix}\right)$
@@
$\mbox{ingest\_y}(s,\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right))=\left(\begin{smallmatrix}(sa+b)&a&(sc+d)&c\\ (se+f)&e&(sg+h)&g\end{smallmatrix}\right)$
for the result of substituting $y\leftarrow s+1/y$.
If one or both inputs are rational, we will eventually ingest infinity, which we define by taking the limit. In most cases
$\mbox{ingest\_x}(\infty,\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right))=\left(\begin{smallmatrix}0&0&a&b\\ 0&0&e&f\end{smallmatrix}\right)$
$\mbox{ingest\_y}(\infty,\left(\begin{smallmatrix}a&b&c&d\\ e&f&g&h\end{smallmatrix}\right))=\left(\begin{smallmatrix}0&a&0&c\\ 0&e&0&g\end{smallmatrix}\right)\ .$
However, this definition of ingest_x is valid only when the numerator and denominator both contain $x$ terms (i.e. only if $a,b$ are not both zero and $e,f$ are not both zero), with an analogous requirement for ingest_y. If the matrix is of the form $\frac{bx+d}{fy+h}$ then the limit as $x\to\infty$ is infinity777with the caveat that the denominator cannot be zero, which occurs if we divide by zero; see section 3.6. and the limit as $y\to\infty$ is zero.
 {input $x,y=[x_{0},x_{1},\cdots],[y_{0},y_{1},\cdots]$}
@@
$M\leftarrow\mbox{ingest\_y}(y_{0},\mbox{ingest\_x}(x_{0},M))${updating $M$ updates variables $a,\cdots,h$}
@@
Figure 2: Preliminary algorithm for arithmetic on two CFs
We now have a preliminary algorithm for arithmetic, which is essentially the original HAKMEM algorithm in more explicit notation. This algorithm will in fact require substantial modification, but we include it as a summary of the key ideas. It is not much different from the single-CF case. Computations of x+y,xy,xyx+y,x-y,xy, and $x/y$ begin with matrices $\left(\begin{smallmatrix}0&1&1&0\\ 0&0&0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}0&1&-1&0\\ 0&0&0&1\end{smallmatrix}\right),\left(\begin{smallmatrix}1&0&0&0\\ 0&0&0&1\end{smallmatrix}\right)$ and $\left(\begin{smallmatrix}0&1&0&0\\ 0&0&1&0\end{smallmatrix}\right)$.
@@
3.1 Failure to Converge
Our preliminary algorithm demonstrates the basic idea of CF arithmetic, but is insufficient because it can fail to converge; we might endlessly ingest terms of both $x$ and $y$ without ever obtaining a bihomographic expression whose floor is known. The simplest case is multiplying $\sqrt{2}=[1,(2)]$ by itself. The first iteration, ingesting 1 from $x$ and from $y$, gives
$z=\sqrt{2}*\sqrt{2}=[z_{0},z_{1},\cdots]=\frac{xy+x+y+1}{xy}\ .$